# Contextuality from Single-State Representations: An Information-Theoretic Principle for Adaptive Intelligence

### v1

@@ VI Implications for Adaptive Intelligence
The results of this work have direct implications for adaptive and intelligent systems operating under internal resource constraints. Many such systems must interact with their environment across multiple contexts while reusing a fixed internal state space, due to limitations on memory, representational capacity, or physical degrees of freedom. Our results show that, under these conditions, contextual dependence cannot be represented classically without incurring an irreducible information-theoretic cost.
From this perspective, phenomena such as context sensitivity, order effects, and apparent violations of classical probabilistic reasoning need not be viewed as anomalies or implementation flaws. Instead, they arise naturally as consequences of representational constraints imposed by single-state reuse. Classical representations fail not because they are insufficiently expressive, but because they require additional contextual information that cannot be accommodated within a fixed internal state space.
@@ VII Discussion and Outlook
We have shown that maintaining a fixed internal state space across multiple interventions entails an irreducible information-theoretic cost for classical probabilistic representations whenever contextual dependence is present. This result reframes contextuality as a representational constraint rather than a peculiarity of specific physical theories. The obstruction we identify is principle-level: it is independent of task details, dynamics, learning rules, or physical implementation.
Several clarifications are in order. First, our results do not imply that classical models are universally inadequate, nor that contextual dependence cannot be handled by classical means in general. Rather, the obstruction arises specifically under the requirement of single-state reuse without context-dependent refinement. Relaxing this requirement—by allowing explicit context encoding, additional memory, or intervention-indexed state spaces—removes the obstruction at the cost of increased representational resources. Our analysis thus makes explicit a trade-off that is often left implicit in classical modeling.
Second, while nonclassical probabilistic frameworks naturally accommodate contextual dependence under single-state constraints, our results do not mandate any particular formalism. Quantum probability appears here as one consistent realization of a broader class of nonclassical representations, distinguished not by physical assumptions but by representational structure. Identifying the minimal mathematical ingredients required for such representations remains an open direction for future work.
Finally, we emphasize that an intervention is not assumed to store its identity within the internal state, nor to leave any persistent record accessible to subsequent interventions. Operationally, interventions are memoryless transformations: once the outcome is produced, no record of the applied intervention is retained in the system. This assumption parallels standard physical idealizations such as projective measurement followed by state update, or controlled operations with subsequent reset. Allowing interventions to leave persistent traces would amount to introducing additional memory resources, thereby explicitly violating the single-state constraint.
Looking ahead, the representational perspective developed here suggests several avenues for further investigation. These include sharpening the relationship between information-theoretic cost and representational complexity, exploring extensions to dynamical and learning settings, and examining how similar obstructions manifest in multi-agent or hierarchical systems. More broadly, our results point toward a unified view of contextuality as a fundamental constraint on representation in bounded adaptive systems, bridging concepts from information theory, probability, and the theory of intelligence.
Related work on information- and memory-cost analyses of simulating contextual correlations has been developed in quantum information theory. Our contribution is not to optimize simulation costs for particular contextuality scenarios, but to isolate a representation-level obstruction under the single-state reuse requirement with interventions-as-operations rather than context labels. In this sense, the present theorem is formulated as a general representational principle, independent of specific inequality families or physical implementations.
@@ Appendix A Proof of Theorem 1
This appendix provides a proof of the information-theoretic obstruction stated in Theorem 1. The proof is representation-theoretic in nature and does not rely on any assumptions about dynamics, learning rules, or physical realization.
Throughout this appendix, we express the abstract shorthand $H>I$ used in the main text in the explicit form
$H(M)\geq I(C;O\mid S),$
where $C$ denotes the intervention, $S$ the internal state, $O$ the observable outcome, and $M$ an auxiliary contextual variable.
@@ A.1 Setup
We consider a classical probabilistic representation with a fixed internal state space $\mathcal{S}$, satisfying the single-state and consistency conditions of Definitions 1–3. Let $S\in\mathcal{S}$ denote the internal state random variable (with realizations $s$), and let $\mathcal{I}=\{I_{1},I_{2},\ldots\}$ denote a set of interventions. Observable outcomes are generated according to $p(o\mid s,I_{k})$.
By assumption, all observable statistics across interventions admit a single underlying classical probability space. Contextual dependence is present when no joint distribution over outcomes exists that is consistent with all observed marginals without reference to the intervention index.
A.2 Contextual Dependence and Incompatible Conditionals
Under the single-state constraint, the same internal state $s$ must support all interventions. Contextual dependence therefore manifests as incompatibility among the families of conditional distributions $\{p(o\mid s,I_{k})\}_{k}$ when viewed as arising from a single underlying random variable $s$.
To reconcile these incompatible conditionals within a classical probability space, additional information specifying how the representation should resolve contextual variation is required. This information cannot be encoded in $s$ itself without introducing intervention-dependent refinement of the state space, which is prohibited by the single-state condition.
@@ A.3 Information-Theoretic Bound
In any classical single-state realization with an auxiliary contextual register $M$, outcomes are generated by a channel $C\to(S,M)\to O$, i.e., $p(o\mid s,m,c)=p(o\mid s,m)$ once $M$ carries all context-dependence not contained in $S$. By the data-processing inequality,
$I(C;O\mid S)\;\leq\;I(C;M\mid S).$ (2)
Moreover, $I(C;M\mid S)\leq H(M)$ holds for any random variables. Hence
$I(C;O\mid S)\;\leq\;H(M).$ (3)
Finally, contextual dependence implies $I(C;O\mid S)>0$ for any single-state classical representation that keeps $S$ context-independent; therefore $H(M)>0$ is necessary, yielding Eq. (1). The bound $H(M)\geq I(C;O\mid S)$ is saturated when $M$ carries precisely the minimal information required to distinguish incompatible conditional structures, for example when $M=f(C)$ in the construction of Sec. IV, and no additional correlations with $S$ or $O$ are introduced.
A.4 Independence from Capacity and Dynamics
Importantly, the bound above does not arise from insufficient capacity of the internal state space. Even if $\mathcal{S}$ is arbitrarily large, the requirement that it remain fixed and reused across interventions prevents the absorption of contextual information into $s$ without violating representation consistency. Likewise, the argument does not depend on any specific dynamics or learning procedures, relying solely on representational constraints.
This completes the proof.

### v2

@@ VI Implications for Adaptive Intelligence
We emphasize that the following discussion is interpretive and does not enter the proof of the main no-go theorem. The results of this work have direct implications for adaptive and intelligent systems operating under internal resource constraints. Many such systems must interact with their environment across multiple contexts while reusing a fixed ontic state space, due to limitations on memory, representational capacity, or physical degrees of freedom. Our results show that, under these conditions, contextual dependence cannot be represented classically without incurring an irreducible information-theoretic cost.
From this perspective, phenomena such as context sensitivity, order effects, and apparent violations of classical probabilistic reasoning need not be viewed as anomalies or implementation flaws. Instead, they arise naturally as consequences of representational constraints imposed by single-state reuse. Classical representations fail not because they are insufficiently expressive, but because they require additional contextual information that cannot be accommodated within a fixed ontic state space.
@@ VII Discussion and Outlook
We have established an information-theoretic no-go theorem for classical ontological models constrained to reuse a single ontic state space across multiple interventions. Specifically, we have shown that contextual dependence necessarily incurs an irreducible contextual information cost quantified by the conditional mutual information $I(C;O\mid\lambda)$. This result identifies contextuality as a fundamental representational constraint arising from limitations on information reuse within classical ontological models.
Our result complements existing no-go theorems such as the Kochen–Specker theorem [9, 12] by identifying an explicit information-theoretic obstruction. Rather than arising from algebraic properties of observables or Hilbert space structure, the obstruction emerges directly from the requirement that a single ontic state space be reused across multiple interventions without contextual refinement. This provides a new perspective on contextuality as an information-theoretic limitation on classical representations.
Nonclassical probabilistic frameworks avoid this obstruction by relaxing the assumption of a single global classical probability space. In quantum theory, a single quantum state $\rho$ is reused across multiple measurement contexts, but measurement outcomes are not assumed to arise from a single underlying classical ontic variable reproducing all measurement statistics simultaneously. This structural feature allows contextual dependence to be represented without introducing auxiliary contextual information external to the quantum state.
Beyond quantum foundations, similar representational constraints arise naturally in adaptive and resource-limited systems. For example, perceptual systems with limited working memory, decision-making agents with bounded internal representations, and physical devices with fixed internal degrees of freedom must often reuse a single ontic state space across multiple interactions. Related constraints have been studied extensively in cognitive science and artificial intelligence under notions such as bounded rationality, resource-limited representation, and working memory limitations [11, 2, 3, 5]. From this perspective, contextuality can be understood more generally as a consequence of representational constraints imposed by ontic state reuse.
Our analysis is representation-theoretic and does not depend on specific physical dynamics or implementation details. The obstruction arises solely from the structural requirement that a single ontic state space be reused across multiple interventions. Relaxing this requirement—for example, by allowing context-dependent ontic state refinement or additional contextual memory—removes the obstruction at the cost of increased representational resources.
Several directions for future work remain. These include extending the analysis to dynamical settings, characterizing minimal contextual information costs for specific contextuality scenarios, and exploring connections between contextuality, information complexity, and representational efficiency. More broadly, the information-theoretic perspective developed here may provide a useful framework for understanding fundamental limits on classical representations in both physical and abstract systems.
In summary, our results establish contextuality as a fundamental information-theoretic obstruction to classical ontological models constrained by ontic state reuse. This provides a new foundation for understanding contextuality as a general representational constraint arising from limitations on classical probabilistic structure.
@@ Appendix A Proof of Theorem 1
This appendix provides a proof of Theorem 1. The proof is representation-theoretic in nature and does not rely on any assumptions about physical dynamics, learning rules, or implementation details. Throughout this appendix, we express the information-theoretic obstruction in explicit form as
$H(M)\;\geq\;I(C;O\mid\lambda),$ (6)
where $C$ denotes the intervention (measurement context), $\lambda\in\Lambda$ the ontic state, $O$ the observable outcome, and $M$ an auxiliary contextual variable not contained in the ontic state.
All entropies and mutual informations are Shannon quantities [10, 6].
@@ A.1 Setup
We consider a classical ontological model with a fixed ontic state space $\Lambda$, satisfying the single-state conditions of Definition 2. The ontic state is represented by a random variable $\lambda\in\Lambda$, distributed according to a preparation distribution $\mu(\lambda)$.
Let $\mathcal{C}$ denote the set of interventions (measurement contexts), with $C\in\mathcal{C}$ a random variable specifying the selected intervention. Observable outcomes are generated according to response functions $\xi(o\mid C,\lambda)$, yielding operational statistics
$p(o\mid C)=\int_{\Lambda}\mu(\lambda)\,\xi(o\mid C,\lambda)\,d\lambda.$ (7)
By assumption, the ontic state space $\Lambda$ is fixed and reused across all interventions, and all observable statistics arise from a single underlying classical probability space over $\Lambda$.
Contextual dependence is present when the observable statistics cannot be represented as arising from a context-independent ontic description alone. Formally, this corresponds to the condition
$I(C;O\mid\lambda)>0.$ (8)
This inequality expresses that knowledge of the intervention $C$ provides additional information about the outcome $O$ beyond what is contained in the ontic state $\lambda$.
A.2 Necessity of an Auxiliary Contextual Variable
Under the single-state constraint, the ontic state space $\Lambda$ is fixed and cannot be refined or indexed by the intervention. Therefore, any contextual dependence that cannot be encoded in $\lambda$ must be mediated through an additional contextual variable $M$.
Operationally, the most general classical ontological model reproducing the statistics can be represented by the probabilistic channel
$C\;\to\;(\lambda,M)\;\to\;O,$ (9)
such that the outcome distribution satisfies
$p(o\mid\lambda,M,C)=p(o\mid\lambda,M).$ (10)
This condition expresses that once the ontic state $\lambda$ and auxiliary contextual variable $M$ are specified, the outcome is independent of the intervention.
@@ A.3 Information-Theoretic Bound
We now derive the lower bound on the contextual information required.
From the channel structure
$C\;\to\;(\lambda,M)\;\to\;O,$ (11)
the data-processing inequality implies
$I(C;O\mid\lambda)\;\leq\;I(C;M\mid\lambda).$ (12)
This inequality expresses that any contextual dependence observed in $O$ must be mediated through the auxiliary contextual variable $M$ once the ontic state $\lambda$ is fixed.
Next, using the general information-theoretic bound
$I(C;M\mid\lambda)\;\leq\;H(M),$ (13)
we obtain
$I(C;O\mid\lambda)\;\leq\;H(M).$ (14)
Rearranging yields the fundamental bound
$H(M)\;\geq\;I(C;O\mid\lambda).$ (15)
Finally, contextual dependence implies
$I(C;O\mid\lambda)>0,$ (16)
since otherwise the outcome would be conditionally independent of the intervention given the ontic state, contradicting contextual dependence.
Therefore,
$H(M)>0,$ (17)
showing that a strictly positive amount of contextual information is required.
A.4 Minimality and Saturation
The bound is tight. Equality
$H(M)=I(C;O\mid\lambda)$ (18)
is achieved when the auxiliary contextual variable $M$ carries precisely the minimal information required to resolve contextual dependence, without introducing additional correlations.
An explicit example achieving saturation is given in Sec. IV, where $M$ is constructed as a deterministic function of the intervention $C$ encoding the minimal contextual information necessary to reproduce incompatible conditional structures.
A.5 Independence from Ontic State Capacity and Dynamics
Importantly, the bound does not arise from limitations on the size or capacity of the ontic state space $\Lambda$. Even if $\Lambda$ is arbitrarily large, the requirement that it remain fixed across interventions prevents contextual information from being absorbed into the ontic state without violating the single-state condition.
Likewise, the argument does not depend on any assumptions about physical dynamics, measurement disturbance, or learning procedures. The obstruction arises purely from representational constraints imposed by ontic state reuse.
A.6 Conclusion of the Proof
We have shown that any classical ontological model constrained to reuse a single ontic state space must introduce an auxiliary contextual variable $M$ satisfying
$H(M)\;\geq\;I(C;O\mid\lambda)\;>\;0.$ (19)
This establishes the information-theoretic no-go theorem stated in Theorem 1.
$\square$